OpenStudy (anonymous):

Solve the equation. Show your work. 8x – 12 = 4x + 24

OpenStudy (anonymous):

x=9

jhonyy9 (jhonyy9):

4x=36 x=36/4 =9

OpenStudy (anonymous):

8x – 12 = 4x + 24 +12 +12 8x =4x +24 +12 4x=36 x=36/4 =9

OpenStudy (anonymous):

try to combine like terms then solve for x

OpenStudy (anonymous):

\[8x-12=4x+24\]\[8x-4x=24+12\]\[4x=36\]\[x=9\]